# Breaking Data Silos: Cross-Domain Learning for Multi-Agent Perception from Independent Private Sources

### 06 Feb 2024 v1

@@ Abstract
The diverse agents in multi-agent perception systems may be from different companies. Each company might use the identical classic neural network architecture based encoder for feature extraction. However, the data source to train the various agents is independent and private in each company, leading to the Distribution Gap of different private data for training distinct agents in multi-agent perception system. The data silos by the above Distribution Gap could result in a significant performance decline in multi-agent perception. In this paper, we thoroughly examine the impact of the distribution gap on existing multi-agent perception systems. To break the data silos, we introduce the Feature Distribution-aware Aggregation (FDA) framework for cross-domain learning to mitigate the above Distribution Gap in multi-agent perception. FDA comprises two key components: Learnable Feature Compensation Module and Distribution-aware Statistical Consistency Module, both aimed at enhancing intermediate features to minimize the distribution gap among multi-agent features. Intensive experiments on the public OPV2V and V2XSet datasets underscore FDA’s effectiveness in point cloud-based 3D object detection, presenting it as an invaluable augmentation to existing multi-agent perception systems.
@@ I Introduction
Existing multi-agent perception methods often operate under a strong assumption: all agents are trained using identical training data. However, this assumption might be not true. The diverse agents in multi-agent perception systems may be from different companies. Each company might use the identical classic neural network architecture based encoder for feature extraction. However, the data source to train the various agents is independent and private in each company as depicted in Fig. 1, leading to the Distribution Gap of different private data for training distinct agents in multi-agent perception system. As an example of Distribution Gaps, let us compare the two widely-used large-scale simulated V2X cooperative perception datasets, namely OPV2V [5] and V2XSet [8]. We find key differences in the following two perspectives.
Agent Number and Type: The presence of diverse agent numbers within the same scenario can yield a different pool of shared visual information. Variations in the quantity and types of agents, e.g., Connected and Automated Vehicles (CAVs) and Infrastructures, can also contribute to disparities in the sharing of intermediate features.
@@
The data silos by the above Distribution Gap could result in a significant performance decline in multi-agent perception. A naive solution for solving this problem might be using federated learning based methods. Federated learning represents a privacy-preserving collaborative machine learning paradigm [9, 10], designed to facilitate the collaborative training of a shared global model among multiple clients by exchanging information pertaining to model parameters. However, in the autonomous driving industry, the data collection and annotation are extremely expensive and with business secrets, so the companies usually reject to share the independent and private data sources due to business competition. As a result, breaking data silos for multi-agent perception is never studied before.
To address the Distribution Gap issue, we introduce the Feature Distribution-aware Aggregation framework for Multi-agent perception system on 3D object detection task, named as FDA. We choose the task of V2X cooperative perception for the point cloud-based 3D object detection as investigation in this paper. Specifically, we proposed two key components: 1) Learnable Feature Compensation Module (LFCM) to generate a residual compensation map for CAV’s intermediate feature considering the large-scale feature information; 2) Distribution-aware Statistical Consistency Module (DSCM) to diminish the distribution gap between CAV feature and ego feature in terms of feature distributions. We conducted extensive experiments on two public V2X perception datasets, namely OPV2V [5] and V2XSet [8], to justify the effectiveness of our proposed method. Our contributions are summarized as follows:
@@
We propose novel Learnable Feature Compensation Module (LFCM) to effectively mitigate discrepancies arising from various agents and Distribution-aware Statistical Consistency Module (DSCM) to diminish the differences between intermediate features extracted from the ego agent and other agents.
@@ II Related Work
V2X-ViT [8] presents a unified ViT architecture for V2X perception which is able to capture the heterogeneous nature in V2X systems.
@@ III-A Overview of the Feature Distribution-aware Aggregation
The Feature Distribution-aware Aggregation framework via the V2V cooperative perception pipeline is illustrated in Fig. 2. Frist, we select an ego vehicle among the CAVs to create a spatial graph that encompasses nearby CAVs (here, we straightforwardly categorize infrastructure as CAV-related.) within the communication range. All the other nearby CAVs will project their own LiDAR data onto the ego vehicle’s coordinate frame, based on both the ego vehicle’s and their own GPS poses. The point clouds from the ego and CAVs are denoted as and , respectively. Each CAV has its own Encoder for LiDAR feature extraction. After feature extraction of each CAV, the ego vehicle receives the neighboring CAV visual features via V2V communication. The intermediate features aggregated from other surrounding CAVs are denoted as , and the ego intermediate features are denoted as . These, along with features from other CAVs, are fed into our proposed LFRM and DSCM modules to reduce feature disparities. After processing by our proposed modules, the intermediate features are fused using the Feature Fusion Network (FFN). Finally, the resulting fused feature maps are fed into a prediction header for 3D bounding-box regression and classification. We formulate the original V2V cooperative perception system for LiDAR-based 3D object detection as
where is the Feature Fusion Network responsible for fusing the features of CAVs and the ego vehicle, and is the prediction header for 3D object detection. Existing cooperative perception methods assume that the encoder weights of and are shared, indicating they are trained with the same distribution data (i.e., ). However, in the real world deployment, , leading to a distribution gap between CAVs and the ego vehicle for most existing models. To address this, we propose a feature distribution-aware aggregation framework, which can serve as a plug-in module in the V2V cooperative perception system. It can be simplified as
@@ III-B Learnable Feature Compensation Module
After encoding by and , we obtain intermediate neural features and from the ego vehicle and other CAVs, respectively. Due to the difference between and based on distinct training data, even when the same point cloud is fed into both encoders, different feature maps are obtained. To mitigate the impact of varying weights in the CAV encoder , we introduce a Learnable Feature Compensation Module (LFCM). The LFCM is an encoder-decoder architecture with skip connections, inspired by the work of [24], and depicted in Fig. 2. The encoder and decoder components of our LFCM employ convolutional layers to capture large-scale spatial features. The shared features from CAVs are fed into the LFCM to predict residual compensation scores for the entire feature set, generating a spatial compensation map of the same size as . Subsequently, is added to , incorporating residual compensation levels into consideration, resulting in the enhanced compensated feature . The computation of the proposed LFCM is formulated as
@@ III-C Distribution-aware Statistical Consistency Module
After processing with our LFCM, we obtain enhanced compensated features for the CAV, denoted as , and the ego intermediate features . To address the distribution gap between and , it is crucial to examine the relationship between statistical differences and feature distributions. Previous studies [25] have established a positive correlation between statistical differences and distribution disparities. In order to minimize the discrepancy in feature distributions between and , we introduce the Maximum Mean Discrepancy (MMD) [26] distance as a metric. Let and represent sets of ego and compensated CAV features, respectively. Our goal is to enhance the compensated CAV features to reduce the distribution gap with ego features in terms of feature distributions by tuning the parameters of . This process can be formulated as follows:
Here, represents the MMD loss between the two sets of intermediate features.
@@ IV-A Experimental Setup
Experimental Details: In our investigation aimed at addressing the distribution gap in LiDAR-based 3D object detection, we make the assumption that all agents employ the same cooperative perception method. However, there is a distinction between the weight of encoders used for the ego vehicle and the CAVs, denoted as . Unlike traditional training strategies where all encoders are trained on the same dataset, we introduce a novel approach. Specifically, we use the OPV2V and V2XSet training sets to train the ego encoder and CAV encoders, respectively. This configuration involves scenarios where the CAV encoders are trained on the OPV2V training set, while the ego encoder is trained on the V2XSet training set, referred to as OPV2V V2XSet. The testing sets of these datasets (i.e., OPV2V and V2XSet) are used to evaluate our cooperative methods. We assess our models under two key settings:
@@
Implementation Details: All cooperative perception models utilize PointPillar [29] as the backbone. We employ the Adam optimizer [30] with an initial learning rate of , gradually decaying it every 10 epochs using a factor of 0.1. The hyperparameters align with those [8]. To simulate real-world scenarios where CAV encoders originate from diverse sources that are challenging to update, we freeze the parameters of CAV encoders during finetuning. All models are trained using two RTX 3090 GPUs. The coefficients for the detection loss () and our MMD loss () are both set to 1.0.
@@ IV-B Quantitative Evaluation
Under the V2XSet OPV2V setting, observable disparities exist between ego and CAVs’ features. Our FDA effectively alleviates this gap, rendering the features more akin to those of the ego.
For V2X-ViT, as depicted in (d) of Fig. 3, its global feature styles of CAVs become more similar to those shown in (b). In the case of CoBEVT, our FDA enhances local details, narrowing the gap with ego’s features, as seen in (f) and (h). This visual evidence underscores the effectiveness of our FDA.
@@ V Conclusions
In this paper, we present pioneering research on the Distribution Gap, stemming from disparate private training data for distinct agents in multi-agent perception systems, resulting in data silos. We analyze the impact of the Distribution Gap on existing cooperative perception methods. To mitigate data silos, we propose a Feature Distribution-aware Aggregation framework, comprising a Learnable Feature Compensation Module and a Distribution-aware Statistical Consistency Module. We empirically assess the FDA framework on two extensive simulated datasets, OPV2V and V2XSet, demonstrating its superior performance in point cloud-based 3D object detection. Our findings emphasize the crucial role of addressing the distribution gap in multi-agent perception systems, contributing to the advancement of cooperative perception strategies with potential implications for autonomous driving.

### 20 Feb 2024 v2

@@ Abstract
The diverse agents in multi-agent perception systems may be from different companies. Each company might use the identical classic neural network architecture based encoder for feature extraction. However, the data source to train the various agents is independent and private in each company, leading to the Distribution Gap of different private data for training distinct agents in multi-agent perception system. The data silos by the above Distribution Gap could result in a significant performance decline in multi-agent perception. In this paper, we thoroughly examine the impact of the distribution gap on existing multi-agent perception systems. To break the data silos, we introduce the Feature Distribution-aware Aggregation (FDA) framework for cross-domain learning to mitigate the above Distribution Gap in multi-agent perception. FDA comprises two key components: Learnable Feature Compensation Module and Distribution-aware Statistical Consistency Module, both aimed at enhancing intermediate features to minimize the distribution gap among multi-agent features. Intensive experiments on the public OPV2V and V2XSet datasets underscore FDA’s effectiveness in point cloud-based 3D object detection, presenting it as an invaluable augmentation to existing multi-agent perception systems. The code is available at https://github.com/jinlong17/BDS-V2V.
@@ I Introduction
Existing multi-agent perception methods often operate under a strong assumption: all agents are trained using identical training data. However, this assumption might be not true. The diverse agents in multi-agent perception systems may be from different companies. Some companies whose perception systems are established on the same popular open platform might use the identical classic neural network architecture-based encoder for feature extraction. However, the data source to train the various agents is independent and private in each company as depicted in Fig. 1, leading to the Distribution Gap of different private data for training distinct agents in multi-agent perception systems. As an example of Distribution Gaps, let us compare the two widely-used large-scale simulated V2X cooperative perception datasets, namely OPV2V [5] and V2XSet [8]. We find key differences in the following two perspectives.
Agent Number and Type: The presence of diverse agent numbers within the same scenario can yield a different pool of shared visual information. Variations in the quantity and types of agents, e.g., Connected and Automated Vehicles (CAVs), and Infrastructures can also contribute to disparities in the sharing of intermediate features.
@@
The data silos by the above Distribution Gap could result in a significant performance decline in multi-agent perception. A naive solution for solving this problem might be using federated learning-based methods. Federated learning represents a privacy-preserving collaborative machine learning paradigm [9, 10], designed to facilitate the collaborative training of a shared global model among multiple clients by exchanging information pertaining to model parameters. However, in the autonomous driving industry, the data collection and annotation are extremely expensive and with business secrets, so the companies usually reject to share independent and private data sources due to business competition and privacy concerns. As a result, breaking data silos for multi-agent perception is never studied before.
To address the Distribution Gap issue, we introduce the Feature Distribution-aware Aggregation framework for Multi-agent perception system on 3D object detection task, named as FDA. We choose the task of V2X cooperative perception for the point cloud-based 3D object detection as an investigation in this paper. Specifically, we proposed two key components: 1) Learnable Feature Compensation Module (LFCM) to generate a residual compensation map for CAV’s intermediate feature considering the large-scale feature information; 2) Distribution-aware Statistical Consistency Module (DSCM) to diminish the distribution gap between CAV feature and ego feature in terms of feature distributions. We conducted extensive experiments on two public V2X perception datasets, namely OPV2V [5] and V2XSet [8], to justify the effectiveness of our proposed method. Our contributions are summarized as follows:
@@
We propose a novel Learnable Feature Compensation Module (LFCM) to effectively mitigate discrepancies arising from various agents and a Distribution-aware Statistical Consistency Module (DSCM) to diminish the differences between intermediate features extracted from the ego agent and other agents.
@@ II Related Work
V2X-ViT [8] presents a unified ViT architecture for V2X perception which is able to capture the heterogeneous nature of V2X systems.
@@ III-A Overview of the Feature Distribution-aware Aggregation
The Feature Distribution-aware Aggregation framework via the V2V cooperative perception pipeline is illustrated in Fig. 2. First, we select an ego vehicle among the CAVs to create a spatial graph that encompasses nearby CAVs within the communication range. Because of the similar sharing capabilities of CAVs and infrastructures, we consider each intelligent infrastructure as a CAV during methodology development in this paper. All the other nearby CAVs will project their own LiDAR data onto the ego vehicle’s coordinate frame, based on both the ego vehicle’s and their own GPS poses. The point clouds from the ego and CAVs are denoted as and , respectively. Each CAV has its own Encoder for LiDAR feature extraction. After feature extraction of each CAV, the ego vehicle receives the neighboring CAV visual features via V2V communication. The intermediate features aggregated from surrounding CAVs are denoted as , and the ego intermediate features are denoted as . These, along with features from other CAVs, are fed into our proposed LFCM and DSCM modules to reduce feature disparities. After processing by our proposed modules, the intermediate features are fused using the Feature Fusion Network (FFN). Finally, the resulting fused feature maps are fed into a prediction header for 3D bounding-box regression and classification. We formulate the original V2V cooperative perception system for LiDAR-based 3D object detection as
where is the Feature Fusion Network responsible for fusing the features of CAVs and the ego vehicle, and is the prediction header for 3D object detection. Existing cooperative perception methods assume that the encoder weights () of and are shared, indicating they are trained with the same distribution data (i.e., ). However, in the real-world deployment, , leading to a distribution gap between CAVs and the ego vehicle for most existing models. To address this, we propose a feature distribution-aware aggregation framework, which can serve as a plug-in module in the V2V cooperative perception system. Then, the Eq. (1) can be modified as
@@ III-B Learnable Feature Compensation Module
After encoding by and , we obtain intermediate neural features and from the ego vehicle and other CAVs, respectively. Due to the difference between and based on distinct training data, even when the same point cloud is fed into both encoders, different feature maps are obtained. To mitigate the impact of varying weights in the encoders, we introduce a Learnable Feature Compensation Module (LFCM). The LFCM is an encoder-decoder architecture with skip connections, inspired by the work of [24], and depicted in Fig. 2. The encoder and decoder components of our LFCM employ convolutional layers to capture large-scale spatial features. The shared features from CAVs are fed into the LFCM to predict residual compensation scores for the entire feature set, generating a spatial compensation map of the same size as . Subsequently, is added to , incorporating residual compensation levels into consideration, resulting in the enhanced compensated feature . The computation of the proposed LFCM is formulated as
@@ III-C Distribution-aware Statistical Consistency Module
After processing with our LFCM, we obtain enhanced compensated features for the CAV, denoted as , and the ego intermediate features . To address the distribution gap between and , it is crucial to examine the relationship between statistical differences and feature distributions. Previous studies [25] have established a positive correlation between statistical differences and distribution disparities. In order to minimize the discrepancy in feature distributions between and , we introduce the Maximum Mean Discrepancy (MMD) [26] distance as a metric. Let and represent the sets of ego and compensated CAV features, respectively. Our goal is to enhance the compensated CAV features to reduce the distribution gap with ego features in terms of feature distributions by tuning the parameters of . This process can be formulated as follow:
where represents the MMD loss between the two sets of intermediate features [26].
@@ IV-A Experimental Setup
Experimental Details: In our investigation aimed at addressing the distribution gap in LiDAR-based 3D object detection, we make the assumption that all agents employ the same cooperative perception method. However, there is a distinction between the weight of encoders used for the ego vehicle and the CAVs, denoted as . Unlike traditional training strategies where all encoders are trained on the same dataset, we introduce a novel approach. Specifically, we use the OPV2V and V2XSet training sets to train the ego encoder and CAV encoders, respectively. This configuration involves scenarios where the CAV encoders are trained on the OPV2V training set, while the ego encoder is trained on the V2XSet training set, defined as OPV2V V2XSet. The testing sets of these datasets (i.e., OPV2V and V2XSet) are used to evaluate our cooperative methods. We assess our models under two key settings:
@@
Implementation Details: All cooperative perception models utilize PointPillar [29] as the backbone. We employ the Adam optimizer [30] with an initial learning rate of , gradually decaying it every 10 epochs using a factor of 0.1. The hyperparameters align with those settings of [8]. To simulate real-world scenarios where CAV encoders originate from diverse sources that are challenging to update, we freeze the parameters of CAV encoders during finetuning. All models are trained using two RTX 3090 GPUs. The coefficients for the detection loss () and our MMD loss () are both set to 1.0.
@@ IV-B Quantitative Evaluation
Under the V2XSet OPV2V setting, observable disparities exist between ego and CAVs’ features. Our FDA effectively alleviates this gap, rendering the features more akin to those of the ego. In the case of CoBEVT, our FDA enhances local details, narrowing the gap with ego’s features, as seen in (b-d). This visual evidence underscores the effectiveness of our FDA.
@@ V Conclusions
In this paper, we present pioneering research on the Distribution Gap, stemming from disparate private training data for distinct agents in multi-agent perception systems, resulting in data silos. We analyze the impact of the Distribution Gap on existing cooperative perception methods. To mitigate data silos, we propose a Feature Distribution-aware Aggregation framework, comprising a Learnable Feature Compensation Module and a Distribution-aware Statistical Consistency Module. We empirically assess the FDA framework on two public and extensive simulated datasets, OPV2V and V2XSet, demonstrating its superior performance in point cloud-based 3D object detection. Our findings emphasize the crucial role of addressing the distribution gap in multi-agent perception systems, contributing to the advancement of cooperative perception strategies with potential implications for autonomous driving.